OpenStudy (anonymous):

x=(2^n+3^n+4^n+5^n)/25 where n is a natural number..... Find the value of n for which x is natural number too.

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

that was from me

OpenStudy (anonymous):

Note that for x to be a natural number, \(2^n +3^n +4^n +5^n\) needs to be a multiple of 25.

OpenStudy (anonymous):

YEP

OpenStudy (anonymous):

its equal to solving\[k=\frac{2^n+3^n+4^n}{25}\]

OpenStudy (anonymous):

U sure

OpenStudy (anonymous):

yeah because 25 is a divisor of 5^n for n>1

OpenStudy (anonymous):

n must be even

OpenStudy (anonymous):

@mukushla is correct. Using modular arithmetic might help: \[2^n mod 25+3^n mod25+4^n mod 25 + 5^n mod 25 =2^nmod25+3^nmod25 +4^nmod25\]

OpenStudy (anonymous):

Really can solve each one individually to see when they are equal to \(0mod25\) and then see where there is an intersection.

OpenStudy (anonymous):

\[2^n+3^n+4^n \equiv 2^{n+1}+1 \ \ \text{mod} \ 5\]

OpenStudy (anonymous):

if n is even

OpenStudy (anonymous):

what math is that @Traxter

OpenStudy (anonymous):

we must have\[ 2^{2m+1}+1 \equiv 0\ \ \text{mod} \ 5\]but\[2^{2m+1}+1 \equiv 2(-1)^m+1\ \ \text{mod} \ 5\]so there is no answer

OpenStudy (anonymous):

thats modular arithmetic

OpenStudy (anonymous):

Yes there is no answer

OpenStudy (anonymous):

what is ur reasoning?

OpenStudy (anonymous):

2^1=2 3^1=3 4^1=4 5^1=5 2^2=4 3^2=9 4^2=16 5^2=25 2^3=8 3^3=27 4^3=64 5^3=125 2^4=6 3^4=81 4^4=256 5^4=625 2^5=12 3^5=163 4^5=1024 2^6=24 3^6=489 4^6=4096 . . .

OpenStudy (anonymous):

See the last digits.....

OpenStudy (anonymous):

So, the last digit will be 4 when n=1 6 when n=2 4 when n=3 8 when n=4 AND this pattern continues

OpenStudy (anonymous):

But to be divisible by 25 the last digit must be 5.......Thus there is no answer for x.

OpenStudy (anonymous):

@mukushla got my method

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yeah :)

OpenStudy (anonymous):

But I didnt get yours

OpenStudy (anonymous):

np ... thats modular arithmetic ..u'll learn it :) do u know know about your courses at high school ?

OpenStudy (anonymous):

I just checked my syllabes...... There is no modular arithmetic

OpenStudy (anonymous):

oops